# Matthew Shu1∗       Nishant Balepur2∗       Shi Feng3∗       Jordan Boyd-Graber2 1

### v2

@@ Abstract
Flashcard schedulers rely on 1) student models to predict the flashcards a student knows; and 2) teaching policies to pick which cards to show next via these predictions. Prior student models, however, just use study data like the student’s past responses, ignoring the text on cards. We propose content-aware scheduling, the first schedulers exploiting flashcard content. To give the first evidence that such schedulers enhance student learning, we build $\text{KAR}^{\text{3}}\text{L}$, a simple but effective content-aware student model employing deep knowledge tracing (DKT), retrieval, and BERT to predict student recall. We train $\text{KAR}^{\text{3}}\text{L}$ by collecting a new dataset of 123,143 study logs on diverse trivia questions. $\text{KAR}^{\text{3}}\text{L}$ bests existing student models in AUC and calibration error. To ensure our improved predictions lead to better student learning, we create a novel delta-based teaching policy to deploy $\text{KAR}^{\text{3}}\text{L}$ online. Based on 32 study paths from 27 users, $\text{KAR}^{\text{3}}\text{L}$ improves learning efficiency over SOTA, showing $\text{KAR}^{\text{3}}\text{L}$’s strength and encouraging researchers to look beyond historical study data to fully capture student abilities.  [^1] Our data and code are available at https://github.com/Pinafore/fact-repetition.
@@ 1 Introduction
We have 27 new users study in our app and test their medium-term learning on two facets: learning accuracy, the number of new facts learned, and recall time, the time taken to recall answers. In 32 six-day studies, $\text{KAR}^{\text{3}}\text{L}+\Delta$ maintains learning accuracy while reducing recall time, showing learning efficiency gains over FSRS (section 7.3). $\text{KAR}^{\text{3}}\text{L}+\Delta$, a baseline content-aware scheduler, bests SOTA, revealing the strength of $\text{KAR}^{\text{3}}\text{L}$, motivating future works to build better content-aware schedulers, and encouraging researchers to look beyond study data to capture student abilities. Our contributions are:
@@ 7.2 Test Mode Setup
Medium-term learning has not been deeply studied, so we test two facets of medium-term learning. First, we define learning accuracy Alshammari (2019) as the gain in accuracy on the post-test from when users first see cards (pre-test). Higher learning accuracy means the user learned more facts.
Learning accuracy measures how many facts a user has learned, but not how well or quickly they learned them. Thus, we define recall time Ericsson (1985) as the mean time needed for users to recall answers—the time between when the card is viewed until when the answer is submitted. We compute recall time on two splits: cards answered correctly; and all cards. Lower recall time on cards answered correctly means the student is more familiar with cards they know, as they recall correct answers faster, while lower recall time on all cards means that the student spends less time studying.
Combining these, we adapt learning efficiency (LE) from Rasch and Schnotz (2009) as the total correct answers (2020*20 ∗ post-test accuracy) divided by recall time on all cards—cards learned per second. Higher LE means the student is studying more efficiently, as they learn more cards in the same time.
Model Pre/Post-Test Acc Recall Corr/All LE
@@
Table 4: Post-test metrics for $\text{KAR}^{\text{3}}\text{L}+\Delta$ vs FSRS users. Best in bold. * means $t$-test significance ($p<0.05$). $\text{KAR}^{\text{3}}\text{L}+\Delta$ users have similar pre-test and post-test accuracies compared to FSRS users but much lower recall time, leading to learning efficiency (LE) improvements.
@@ 7.3 User Study Results
We collect 32 six-day study sessions from 27 students with FSRS and $\text{KAR}^{\text{3}}\text{L}+\Delta$ (Table 4). Both schedulers help users achieve similar mastery in test mode, more than doubling pre-test to post-test accuracy ($0.42$ to $0.86$). While learning accuracy is similar, $\text{KAR}^{\text{3}}\text{L}+\Delta$ has lower post-test recall time on cards answered correctly, suggesting that $\text{KAR}^{\text{3}}\text{L}$ users are more familiar with the cards answered correctly. Finally, $\text{KAR}^{\text{3}}\text{L}+\Delta$ users maintain post-test accuracy while reducing recall time on all cards studied. As a result, $\text{KAR}^{\text{3}}\text{L}+\Delta$ users obtain higher learning efficiency (LE) than FSRS users, as users studying with $\text{KAR}^{\text{3}}\text{L}$ learn a similar number of facts with less time taken while studying ($2.74$ vs $2.58$ flashcards learned per second).
Combining offline (Table 1) and online (Table 4) results, $\text{KAR}^{\text{3}}\text{L}+\Delta$: 1) produces more accurate and calibrated recall predictions than FSRS on seen cards; 2) can predict recall on unseen cards, unlike FSRS; 3) matches FSRS in enhancing learning accuracy; and 4) enables users to recall answers faster than FSRS, increasing learning efficiency. Holistically, $\text{KAR}^{\text{3}}\text{L}+\Delta$ bests FSRS. Since $\text{KAR}^{\text{3}}\text{L}+\Delta$ is a baseline content-aware scheduler and it already rivals SOTA, content-aware scheduling is a promising paradigm, which we hope will motivate works to build better content-aware schedulers and look beyond study data to fully capture student abilities.
@@ 8 Conclusion
We introduce and successfully implement the first content-aware scheduler with $\text{KAR}^{\text{3}}\text{L}+\Delta$, a simple but effective model using DKT, BERT, retrieval, and a novel delta-based teaching policy. Our offline evaluation on a newly-collected dataset shows $\text{KAR}^{\text{3}}\text{L}$ provides accurate and well-calibrated recall predictions, while our online evaluation reveals $\text{KAR}^{\text{3}}\text{L}$ improves learning efficiency over SOTA. Thus, we give the first evidence that content-aware schedulers can be used to improve student learning.
Content-aware scheduling enhances personalization in learning tools, as models can infer student knowledge gaps through semantic inferences. Given this strength, we hope future works extend content to modalities beyond text, such as images or audio. In online evaluation, $\text{KAR}^{\text{3}}\text{L}$ surpasses FSRS in learning efficiency, but there are still many facets of learning that can be measured with online studies; these learning metrics could not only be used for evaluation but also training signals. In all, we show the strength of content-aware scheduling, which we hope will motivate works to look beyond single study items and also model relations between study items to fully capture student abilities.
@@ Acknowledgements
We would like to thank members of the CLIP lab at the University of Maryland and external collaborators for their discussions of this work, including Jarrett Ye, Jack Wang, David Martinez, Dayeon Ki, Yu Hou, Yoonjoo Lee, Hyojung Han, Haozhe An, Erik Harpstead, Paulo Carvalho, Rachel Rudinger, Pedro Rodriguez, Denis Peskov, Roger Craig, and Wenyan Li. We additionally thank Nathaniel Klein for helping name $\text{KAR}^{\text{3}}\text{L}$. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber) and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Cloud computing resources were made possible by a gift from Adobe Research.

### v3

@@ Abstract
Flashcard schedulers rely on 1) student models to predict the flashcards a student knows; and 2) teaching policies to pick which cards to show next via these predictions. Prior student models, however, just use study data like the student’s past responses, ignoring the text on cards. We propose content-aware scheduling, the first schedulers exploiting flashcard content. To give the first evidence that such schedulers enhance student learning, we build $\text{KAR}^{\text{3}}\text{L}$, a simple but effective content-aware student model employing deep knowledge tracing (DKT), retrieval, and BERT to predict student recall. We train $\text{KAR}^{\text{3}}\text{L}$ by collecting a new dataset of 123,143 study logs on diverse trivia questions. $\text{KAR}^{\text{3}}\text{L}$ bests existing student models in AUC and calibration error. To ensure our improved predictions lead to better student learning, we create a novel delta-based teaching policy to deploy $\text{KAR}^{\text{3}}\text{L}$ online. Based on 32 study paths from 27 users, $\text{KAR}^{\text{3}}\text{L}$ improves testing throughput over SOTA, showing $\text{KAR}^{\text{3}}\text{L}$’s strength and encouraging researchers to look beyond historical study data to fully capture student abilities.  [^1] Our data and code are available at https://github.com/Pinafore/fact-repetition.
@@ 1 Introduction
We have 27 new users study in our app and test their medium-term learning : response accuracy, the number of new flashcards learned, and response time, the time taken to recall answers. In 32 six-day studies, $\text{KAR}^{\text{3}}\text{L}+\Delta$ maintains learning accuracy while reducing recall time, showing testing throughput gains over FSRS (section 7.3). $\text{KAR}^{\text{3}}\text{L}+\Delta$, a baseline content-aware scheduler, bests SOTA, revealing the strength of $\text{KAR}^{\text{3}}\text{L}$, motivating future works to build better content-aware schedulers, and encouraging researchers to look beyond study data to capture student abilities. Our contributions are:
@@ 7.2 Test Mode Setup
Medium-term learning has not been deeply studied, so we use two measures of memory strength that have been explored in prior research: response accuracy and response time MacLeod and Nelson (1984); Maris and Van der Maas (2012). We define response accuracy Alshammari (2019) as the gain in accuracy on the post-test from when users first see cards (pre-test). Higher response accuracy means that the user has learned more flashcards.
Response accuracy captures how many facts a user has learned, but not how well or quickly they can recall these facts. Thus, we define response time as the mean time needed for users to recall flashcard answers Ericsson (1985)—the time from when the card is viewed until when the answer is submitted. We calculate response time on two splits: flashcards answered correctly; and all flashcards studied. Lower response time on cards answered correctly means the user is more familiar with cards they know, since they recall correct answers faster, while lower response time on all cards means that the user completes the post-test faster.
Model Pre/Post-Test Acc Response Corr/All TTP
@@
Table 4: Post-test metrics for $\text{KAR}^{\text{3}}\text{L}+\Delta$ vs FSRS users. Best in bold. * means $t$-test significance ($p<0.05$). $\text{KAR}^{\text{3}}\text{L}+\Delta$ users have similar pre-test and post-test accuracy versus FSRS users but much lower post-test response time, enhancing testing throughput (TTP).
We combine response accuracy and time into a single post-test performance metric Chignell et al. (2014) and define testing throughput (TTP) as the average total correct answers (2020\;*20 ∗ post-test accuracy) divided by the average response time on all cards—cards answered correctly per second spent on the post-test. Higher TTP indicates that the student has a deeper understanding of the test mode flashcards, as they can correctly answer the same number of flashcards while taking less time.
@@ 7.3 User Study Results
We collect 32 six-day study sessions from 27 students with FSRS and $\text{KAR}^{\text{3}}\text{L}+\Delta$ (Table 4). Both schedulers help users achieve similar mastery in test mode, more than doubling pre-test to post-test accuracy ($0.42$ to $0.86$). While response accuracy is similar, $\text{KAR}^{\text{3}}\text{L}+\Delta$ has lower post-test response time on cards answered correctly, suggesting that $\text{KAR}^{\text{3}}\text{L}$ users are more familiar with the cards answered correctly. Finally, $\text{KAR}^{\text{3}}\text{L}+\Delta$ users maintain post-test accuracy while reducing recall time on all cards studied. As a result, $\text{KAR}^{\text{3}}\text{L}+\Delta$ users obtain higher testing throughput (TTP) than FSRS users, as users studying with $\text{KAR}^{\text{3}}\text{L}$ answer a similar number of facts correctly in less time ($2.74$ vs $2.58$ flashcards answered correctly per second).
Combining offline (Table 1) and online (Table 4) results, $\text{KAR}^{\text{3}}\text{L}+\Delta$: 1) produces more accurate and calibrated recall predictions than FSRS on seen cards; 2) can predict recall on unseen cards, unlike FSRS; 3) matches FSRS in enhancing response accuracy; and 4) enables users to recall answers faster than FSRS, increasing testing throughput. Holistically, $\text{KAR}^{\text{3}}\text{L}+\Delta$ bests FSRS. Since $\text{KAR}^{\text{3}}\text{L}+\Delta$ is a baseline content-aware scheduler and it already rivals SOTA, content-aware scheduling is a promising paradigm, which we hope will motivate works to build better content-aware schedulers and look beyond study data to fully capture student abilities.
@@ 8 Conclusion
We introduce and successfully implement the first content-aware scheduler with $\text{KAR}^{\text{3}}\text{L}+\Delta$, a simple but effective model using DKT, BERT, retrieval, and a novel delta-based teaching policy. Our offline evaluation on a newly-collected dataset shows $\text{KAR}^{\text{3}}\text{L}$ provides accurate and well-calibrated recall predictions, while our online evaluation reveals $\text{KAR}^{\text{3}}\text{L}$ improves testing throughput over SOTA. Thus, we give the first evidence that content-aware schedulers can be used to improve student learning.
Content-aware scheduling enhances personalization in learning tools, as models can infer student knowledge gaps through semantic inferences. Given this strength, we hope future works extend content to modalities beyond text, such as images or audio. In online evaluation, $\text{KAR}^{\text{3}}\text{L}$ surpasses FSRS in testing throughput, but there are still many facets of learning that can be measured with online studies; these learning metrics could not only be used for evaluation but also training signals. In all, we show the strength of content-aware scheduling, which we hope will motivate works to look beyond single study items and also model relations between study items to fully capture student abilities.
@@ Acknowledgements
We would like to thank members of the CLIP lab at the University of Maryland and external collaborators for their discussions of this work, including Jarrett Ye, Jack Wang, David Martinez, Dayeon Ki, Yu Hou, Yoonjoo Lee, Hyojung Han, Haozhe An, Erik Harpstead, Paulo Carvalho, Rachel Rudinger, Pedro Rodriguez, Denis Peskoff, Roger Craig, Sydney Montoya, and Wenyan Li. We additionally thank Nathaniel Klein for helping name $\text{KAR}^{\text{3}}\text{L}$. This material is based upon work supported by the National Science Foundation under Grant No. iis-2403436 (Boyd-Graber) and dge-2236417 (Balepur). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Cloud computing resources were made possible by a gift from Adobe Research.